OpenStudy (writerforever18):

Deon invested his savings in two investment funds. The amount he invested in Fund A was 4 times as much as the amount he invested in Fund B. Fund A returned a 5% profit and Fund B returned a 7% profit. How much did he invest in Fund B, if the total profit from the two funds together was $1620 ?

OpenStudy (mathmale):

Please begin by choosing letters to represent your unknown quantities. You could choose x to represent the amount invested in Fund A and y to represent that invested in Fund B. How are x and y related to one another? Reread the question for the answer to this.

OpenStudy (writerforever18):

How would I put the problem together?

OpenStudy (writerforever18):

I understand what the overall problem is asking me, but I don't understand how to know what order to put the equation in in order to solve it. Could you please help me with that? Thank you!

OpenStudy (writerforever18):

I understand what the overall problem is asking me, but I don't understand how to know what order to put the equation in in order to solve it. Could you please help me with that? Thank you!

OpenStudy (mathmale):

Your math course is likely Algebra. In Algebra we very frequently represent unknown quantities with letters of the alphabet. I spelled this out for you earlier: "Please begin by choosing letters to represent your unknown quantities. You could choose x to represent the amount invested in Fund A and y to represent that invested in Fund B. How are x and y related to one another? Reread the question for the answer to this." Let x be the amount invested in Fund A and y be the amount inv. in Fund B. In which Fund did this guy invest more money? A or B? Based upon that, write a single equation that shows how x is related to y. Please share your result here.

OpenStudy (writerforever18):

ok one minute, I'll try to put this together.

OpenStudy (writerforever18):

Do you think the equation would be 4(0.07)+0.05x=1620?

OpenStudy (mathmale):

I do appreciate your taking the initiative here, but would still like for you to answer my question. How much $ goes into Fund A? How much into B? I asked you to write an equation showing how x and y are related. In other words, you need to express one of these in terms of the other.

OpenStudy (writerforever18):

Oh, ok sorry, It says in the problem that Fund A had a return profit of 5% and Fund B has a return profit of 7%. The total profit from both funds= $1620

OpenStudy (writerforever18):

I'm not completely sure how to write an equation showing how x and y are related.

OpenStudy (mathmale):

"The amount he invested in Fund A was 4 times as much as the amount he invested in Fund B." Since the amount invested in Fund A was x, this sentence translates into the equation x=4y.

OpenStudy (writerforever18):

Ok, so I think in that case, I think the equation would be 4x(0.07)=1620(0.05) Is that right?

OpenStudy (writerforever18):

I mean 4(0.07+x)=1620(0.05)

OpenStudy (mathmale):

Note that x and y are DOLLAR AMOUNTS, and that we don't yet have any clue regarding how many dollars are involved. However, we are told that "Fund A returned a 5% profit and Fund B returned a 7% profit," and that the total profit was $1620. Can you now write an equation showing this fact? This equation will involve solely PROFIT, although x and y (dollar amounts) are included in the equation.

OpenStudy (writerforever18):

Would the equation be 0.05x+0.07y=1620?

OpenStudy (writerforever18):

But then what about the 4 times as much as the amount invested in Fund B? Wouldn't that be 4x?

OpenStudy (mathmale):

Nice job on that second equation. We say here, "We have 2 equations in 2 unknown quantities." Since x=4y, why not substitute 4y for x in your 2nd equation, and then find y? Please show your work.

OpenStudy (writerforever18):

So 0.05(4x)+0.07y=1620?

OpenStudy (mathmale):

But wasn't our first equation "x=4y?" Leave the 0.07y alone; substitute 4y for x in 0.05(x).

OpenStudy (writerforever18):

So just 0.05(4x)=1620? Sorry I'm a tiny bit confused as to what to do with the 0.07 now.

OpenStudy (mathmale):

Sorry, my computer crashed and had to be rebooted. You had 0.05x + 0.07y =$1620. But x=4y. Subst. 4y for x: 0.05(4y) + 0.07y = $1620 => 0.20y + 0.07y = $1620. Solve for y. Next, solve for x, using the y value you've just found.

OpenStudy (writerforever18):

That's ok, Alright that makes more sense. I'll do that and then tell you what answer I get to see if it's correct.

OpenStudy (mathmale):

Great. Very pleased with the way you've stuck with this problem solution. Way to go.

OpenStudy (writerforever18):

I got y=6,000

OpenStudy (mathmale):

So, what would x be? Remember, we have an equation relating x and y already.

OpenStudy (writerforever18):

So then you would just put in what you get for y into the y in the original problem right?

OpenStudy (writerforever18):

0.05x+0.07(6,000)=1620?

OpenStudy (writerforever18):

It doesn't seem right to me, but I got x=24,000

OpenStudy (mathmale):

Check your work. x = 4y. If x= 4y, and y = $6000, x is $24000. Next, determine wehther or not 0.05(24000) + 0.07(6000) = 1620.

OpenStudy (writerforever18):

I typed in the answer I got for Fund B, and it's correct! They were only looking for how much was invested in Fund B.

OpenStudy (mathmale):

And so you chose y=$6,000? Super.

OpenStudy (writerforever18):

It said it was correct, thank you so much! :)

OpenStudy (writerforever18):

But now there's a new problem, would you still be able to help me please?

OpenStudy (mathmale):

Please post your next problem separately from this one. Thx.

OpenStudy (writerforever18):

Ok, Thank you!